10 fluorescent bulbs rated 36W each, and 50 LED lights rated 7W each are continuously tested for 24 hours straight. Determine the kWh consumption of the overall test
1
Expert's answer
2021-09-13T12:05:10-0400

Gives

Number of bulbs=10

Number of LED=50

"P_1=36W\\P_2=7W\\t=24hr"

"P_1=\frac{E_1}{t}\\E_1=P_1\times t_1"

"E_1=36\times24=864Wh"


"E_2=P_2\times t_2=7\times 24=168Wh"

Now n=10 bulb

"P_1=\frac{E_1}{t}\\E_1=n\times P_1\times t_1"


"E_1=10\times36\times24=8640Wh=8.64KWh"

Now n=50 LED

"E_2=n\times P_2\times t_2=50\times7\times24=8400Wh=8.4KWh"

Overall energy test=8.64+8.4=17.04KWh

E=17.04KWh